OpenStudy (anonymous):

The radius of one sphere is twice as great as the radius of a second sphere. Find the ratio of their volumes.

OpenStudy (anonymous):

Volume of a sphere:\[V= {4\over3}\pi r^3\] Let's call your two radii r_1 and r_2, where\[r_1 = 2 r_2\]Then we can quickly see:\[V_1 = {4\over3} \pi {r_1}^3\] and \[V_2 = {4\over3} \pi {r_2}^3\]Substituting in the fact that r_1 is double r_2, we find\[V_1 = {4\over3} \pi {(2r_2)}^3\]The ratio of volumes, if you cancel out the 4/3 and pi's, is:\[{V_1 \over V_2} ={ {(2r_2)}^3 \over {r_2}^3} = 2^3 = 8\]

OpenStudy (anonymous):

The ratio is 8? That doesn't make sense.

OpenStudy (anonymous):

Why not? The units make sense.

OpenStudy (anonymous):

This basically means that the first sphere is 8 times as large as the second sphere.

OpenStudy (anonymous):

Oh!! That's what I didn't get! Thanks!